# Robust automatic brain vessel segmentation in 3D CTA scans using dynamic 4D-CTA data

### v1

@@ 4 Discussion
We proposed an automatic data pre-processing method that utilizes dynamic CTA images acquired at multiple time points. This approach not only reduces annotation time but also enhances the diversity of the training data at no additional cost. Using this protocol, we created DynaVessel, a CTA intracranial vessel segmentation dataset with artery and vein labels comprising 110 images from 27 patients, covering a wide range of acquisition contrasts. Furthermore, we annotated a large evaluation set of 165 images from 14 patients with fine-grained semantic labels for 20 arteries, and we evaluate the impact of training $M_{DV}$ (a state-of-the-art nnUNet) on our data using modified metrics to emphasize sensitivity across different vessel segments. Overall, our evaluation pipeline provides detailed insight on the model’s ability to segment arteries of varying sizes.
@@
Across all vascular regions, $M_{DV}$ achieved superior performance compared to models trained on other datasets of similar or much larger sizes. This performance gap was especially pronounced in veins, where $M_{DV}$ achieved a modified DC of 0.75 to 0.95 depending on the dataset, compared to under 0.01 mDC for the other models. The topology Sensitivity and average directed Hausdorff Distance tell a similar story, with $M_{DV}$ outperforming the other models by a significant margin. Although the performance gap was smaller for arteries, $M_{DV}$ still came on top, with modified DC scores of 0.75 to 0.86 compared to scores ranging from 0.07 to 0.45 for the other models. Overall, our work constitutes a significant step toward the automatic segmentation of all brain vascular structures in CTA scans, with broad applicability as a primary or auxiliary input for the assessment of cerebrovascular pathologies.
Although DynaVessel imbues $M_{DV}$ with robust artery and vein segmentation performance, we note that our fine-grained semantic segmentation evaluations only serve to provide an upper-bound approximation of the performance of models trained using DynaVessel. Future work should address the semantic segmentation scenario by expanding the annotations in the training data to include all 20 artery (or more) sub-classes, establishing guidelines to annotate venous structures in a fine-grained manner, and designing deep learning architectures suited for such a complex task.

### v2

@@ 4 Discussion
We proposed an automatic data pre-processing method that utilizes dynamic CTA images acquired at multiple time points. This approach not only reduces annotation time but also enhances the diversity of the training data at no additional cost. Using this protocol, we created DynaVessel, a CTA intracranial vessel segmentation dataset with artery and vein labels comprising 110 images from 29 patients, covering a wide range of acquisition contrasts. Furthermore, we annotated a large evaluation set of 165 images from 14 patients with fine-grained semantic labels for veins and 20 arteries, and we evaluate the impact of training $M_{DV}$ (a state-of-the-art nnUNet) on our data using modified metrics to emphasize sensitivity across different vessel segments. Overall, our evaluation pipeline provides detailed insight on the model’s ability to segment arteries of varying sizes and across multiple phases.
The $M_{DV}$ performance trends across dynamic CTA phases highlight the dependence of deep learning–based vessel segmentation on image contrast. In our experiments training on a single phase, segmentation performance degrades severely during mismatched phases—such as arterial segmentation in late venous images or venous segmentation in early arterial images—where reduced contrast conspicuity introduces ambiguity. Notably, the difference in HU distributions between the training data and phase-specific test data correlates strongly with performance; for instance, $M_{ven}$ is trained on images with a mean vein HU of 485, resulting in a delta of 87 HU with respect to the mean vein HU of the venous phase ($mDC=0.81$) and a delta of 385 HU with respect to the mean HU of the early arterial phase ($mDC=0.01$). Together, these findings suggest that models trained on single-phase data leverage contrast-driven vessel appearance rather than anatomical priors. On the other hand, models trained on multi-phase data suffer far less performance degradation, although they are still sensitive to changes in contrast. Indeed, the artery segmentation performance of $M_{DV}$ peaks during the arterial phase, when arterial enhancement is maximal and venous structures remain relatively unopacified, whereas vein segmentation achieves its highest accuracy during the venous phase, when venous enhancement dominates. Moreover, $M_{DV}$ still requires the presence of some contrast agent to correctly segment brain vessels, as shown by its relatively poor performance during the early arterial phase.
@@
Regardless, when evaluated across all vascular regions, $M_{DV}$ achieved superior performance compared to models trained on other datasets of similar or much larger sizes. This performance gap was especially pronounced in veins, where $M_{DV}$ achieved a modified DC of 0.75 to 0.95 depending on the dataset, compared to under 0.01 mDC for the other models. The topology Sensitivity and average directed Hausdorff Distance tell a similar story, with $M_{DV}$ outperforming the other models by a significant margin. Although the performance gap was smaller for arteries, $M_{DV}$ still came on top, with modified DC scores of 0.75 to 0.86 compared to scores ranging from 0.07 to 0.45 for the other models. Overall, our work constitutes a significant step toward the automatic segmentation of all brain vascular structures in CTA scans, with broad applicability as a primary or auxiliary input for the assessment of cerebrovascular pathologies. Nevertheless, despite the fact DynaVessel imbues $M_{DV}$ with robust vessel segmentation performance, we note that our fine-grained semantic segmentation evaluations only serve to provide an upper-bound approximation of the performance of models trained using DynaVessel. Future work should address the semantic segmentation scenario by expanding the annotations in the training data to include all 20 artery (or more) sub-classes, establishing guidelines to annotate venous structures in a fine-grained manner, and designing deep learning architectures suited for such a complex task.
@@ 4.0.1 Acknowledgements
Supplementary material
Figure S1: Visualization of the semantic labels for the DynaVessel test data. Created using 3D Slicer.
(a) $X_{a}$
(b) $S_{a}$
(c) $S_{a}^{*}$
(d) $S_{v}$
(e) $S_{v}^{*}$
Figure S2: Axial images of arterial phase CTA ($X_{a}$), bone subtracted arterial and venous phase ($S_{a}$ and $S_{v}$), and the vessel separated arterial and venous phase ($S_{a}^{*}$ and $S_{v}^{*}$) volumes from a sample patient.